Paul is now 24 years old and his cousin Mon is 16 years old. How many years ago was paul thrice as old as Mon?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.
Paul is now 24 years old and his cousin Mon is 16 years old.
How many years ago was paul thrice as old as Mon?
~~~~~~~~~~~~~~~~

x years ago

    24 - x = 3*(16-x)


Simplify and find x

    24 - x = 48 - 3x

    3x - x = 48 - 24

       2x  =    24

        x  =    24/2 = 12.


ANSWER.  12 years ago.

Solved.



Answer by greenestamps(13200)



(1) informally....

The difference in their ages is 24-16 = 8 years. So when Paul was 3 times as old as Mon, their ages were 4 and 12. That was 24-12 = 12 years ago.

ANSWER: 12 years

(2) algebraically (setting up the problem differently than the other tutor)....

x = Mon's age when Paul was 3 times as old
3x = Paul's age then

The difference in their ages then was 3x-x = 2x.

But the difference in their ages is 8 years:
2x = 8
x = 4

Mon was 4 years old when Paul was 3 times as old; that was 16-4 = 12 years ago.

ANSWER: 12 years